OpenStudy (hannah.petrova):

Change the fraction 13/20 to a decimal.

OpenStudy (alexandervonhumboldt2):

100/20=5. multiply the whole fraction by 5/5: \[\frac{ 13 }{ 20 }*\frac{ 5 }{ 5 }\] \[\frac{ 65 }{ 100 }=0.65\]

OpenStudy (hannah.petrova):

oh thanks